Obtain the solution of the differential equation


(x^2)y" - 2xy' + 2y = (log (base e) x) ^2, (x>0),


using the method of undetermined coefficients.

1
Expert's answer
2022-03-29T07:26:36-0400

"x^2y''-2xy'+2y=\\ln ^2x\\\\Homogeneous\\,\\,equation:\\\\x^2y''-2xy'+2y=0\\\\y=zx\\\\y'=z+z'x\\\\y''=2z'+z''x\\\\2z'x^2+z''x^3-2zx-2z'x^2+2zx=0\\\\z''x^3=0\\\\z''=0\\\\z=C_1+C_2x\\\\y=C_1x+C_2x^2\\\\The\\,\\,particular\\,\\,solution\\,\\,of\\,\\,the\\,\\,inhomogeneous\\,\\,equation\\,\\,\\\\y=A\\ln x+B\\ln ^2x+C\\\\y'=\\frac{A}{x}+2B\\frac{\\ln x}{x}\\\\y''=-\\frac{A}{x^2}+\\frac{2B}{x^2}-\\frac{2B\\ln x}{x^2}\\\\-A+2B-2B\\ln x-2A-2B\\ln x+2A\\ln x+2B\\ln ^2x+2C=\\ln ^2x\\\\\\left\\{ \\begin{array}{c}\t-3A+2B+2C=0\\\\\t-4B+2A=0\\\\\t2B=1\\\\\\end{array} \\right. \\Rightarrow \\left\\{ \\begin{array}{c}\tA=\\frac{3}{2}\\\\\tB=\\frac{1}{2}\\\\\tC=\\frac{7}{4}\\\\\\end{array} \\right. \\\\y=C_1x+C_2x^2+\\frac{3}{2}\\ln x+\\frac{1}{2}\\ln ^2x+\\frac{7}{4}"